Two boys with two-way radios that have a range of 2 miles leave camp and walk in opposite directions. If one boy walks 3mph and the other walks 4mph, how long will it take before they lose radio contact?
Answer by addingup(3677)

3+4 = 7 they are moving away from each other at a combined speed of 7mph
2/7 = 0.285714 of an hour
An hour has 60 minutes, so:
0.285714 * 60 = 17 minutes
(you can work it further, you'd come up with 17 minutes and 8 secords)